OpenStudy (anonymous):

Divide. Express your answer in lowest terms. 1/8/ 4 =

OpenStudy (anonymous):

Well, you can divide 1/8 first :P

OpenStudy (anonymous):

this is ame as (1/8) + (4)

OpenStudy (anonymous):

4.125

OpenStudy (anonymous):

or (1/8) /(4/1) and the rul there is "invert (turn upside down) and multiply" so for example 3/4 x 2/3 becomes 3/4 x 3/2

OpenStudy (anonymous):

\[\frac{ 1 }{ 8} \div 4\]

OpenStudy (anonymous):

thats the equation if u got confused

OpenStudy (anonymous):

@BPDlkeme234 @School_is_hard

OpenStudy (anonymous):

whoever answers gets a fan and a medal

OpenStudy (anonymous):

remeber the rule (1/8)/(4/1) = (1/8) x (1/4)

OpenStudy (anonymous):

now rules for multipication of fractions like (1/8) x (1/4) = multiply top by top and bottom by bottom..so what is the answer?

OpenStudy (anonymous):

1/8 divided by 4 you turn 4 into a fraction 4/1 then multiply 1/8 times 4/1 = 4/8 simplify divide by 2 equals 2/4 and you can simplify again divide by 2 equals 1/2

OpenStudy (anonymous):

i meant 1/4

OpenStudy (anonymous):

it would be, 1/8 divided by 4 you turn 4 into a fraction 4/1 then multiply 1/8 times 1/4 = 4/8 simplify divide by 2 equals 2/4 and you can simplify again divide by 2 equals 1/2

OpenStudy (anonymous):

the top two (known as numerators) are 1 x 1, yes?

OpenStudy (anonymous):

1/8 times 1/4 doesnt = 4/8?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then simplify to 1/2

OpenStudy (anonymous):

to multiply two fractions, mu.tiply the top by the top and the bottom by the bottom

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

@school_is_hard is wrong, listen to BPD :)

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

well what is 1/8 x 1/4?

OpenStudy (anonymous):

oh I am, sorry

OpenStudy (anonymous):

0.03125

OpenStudy (anonymous):

thx @swagmaster47

OpenStudy (anonymous):

1/32

OpenStudy (anonymous):

oh yeah, it would be 2/32 or 1/16...right?

OpenStudy (anonymous):

Okay lets try something simpler: what is (1/4) x (1/4)?

OpenStudy (anonymous):

1/16

OpenStudy (anonymous):

remember multiply the top by the top i.e. 1 x 1, and the bottom by the bottom. i.e 4 x 4 = 1/16

OpenStudy (anonymous):



OpenStudy (anonymous):

so answer is 1/6

OpenStudy (anonymous):

thanks for the diagram @swagmaster47

OpenStudy (anonymous):

are you with me @School_is_hard ?

OpenStudy (anonymous):

so its 1/6

OpenStudy (anonymous):

@BPDlkeme234

OpenStudy (anonymous):

@School_is_hard

OpenStudy (anonymous):

seem to have lost @School_is_hard

OpenStudy (anonymous):

thankyou @THEUSER822 and @swagmaster47